# Learning Source Biases: Multi-sourced Misspecifications and Consequences

### v2

@@ A.4 Proofs of results that characterize KL minimizers
$=-\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\sum_{i=1}^{M}\Omega_{i}^{-i}\Delta_{i}=-\overline{\Delta}.\qed$
Proof of Theorem 2.
Part (i): The KL miminization problem considered in Case (III) of HKS (recall Appendix A.1) is
where $\underline{\underline{\lambda}}$ can be any positive number smaller than $1\wedge\lambda_{\min}(\Sigma^{*})$. The solution $(\hat{f},\widehat{\Sigma})$ to this problem is shown to satisfy two properties: First, $\Delta f_{i}\coloneqq\hat{f}_{i}-f_{i}^{*}$ is the same as when the DM knows the true covariance matrix $\Sigma^{*}$ from the outset. Second, $\widehat{\Sigma}=\Sigma^{*}+(\Gamma\Delta f)(\Gamma\Delta f)^{\top}$. The assumption that only a single entry of $f$ is misspecified plays no role in this result. This can be verified by replicating the proof of HKS line by line.
Letting $f=b$ and $\Gamma=\mathbb{I}_{N\times N}$ in the above argument leads us to conclude that in the environment outlined in Section 4.3, the solution to the KL minimization problem would be $\Delta=\hat{b}_{i}-b_{i}^{*}$ and $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$, had our DM been able to choose any SPD matrix whose minimum eigenvalue is bounded below by $\underline{\underline{\lambda}}$. The additional constraint we impose is that the covariance matrix must be of the form $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$, where $\Omega$ is PSD with its minimum eigenvalue bounded below by $\underline{\lambda}<1\wedge\lambda_{\min}(\Omega^{*})$. We are left to verify that the solution to the unconstrained problem satisfies this additional constraint, and so must also solve the constrained optimization problem.
To verify, note that the solution to the unconstrained problem: $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$, is attained by $\widehat{\Omega}=\Omega^{*}+\Delta\Delta^{\top}$. $\widehat{\Omega}$ is clearly PSD, and $\lambda_{\min}(\widehat{\Omega})\geq\lambda_{\min}(\Omega^{*})>\underline{\lambda}$ by Lemma 5. This completes the proof of Part (i) of the theorem.
Part (ii) of the theorem follows from the decomposition exercise conducted in Section 4.3. Part (iii) follows from straightforward algebra. ∎
Proof of Proposition 4.
Consider first the setting examined in Section 3. There, $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\overline{\Delta}$, and so
@@
Consider next the setting examined in Section 4.3. There,
@@
$=-\underbrace{\mathbb{E}_{(b^{*},\Sigma^{*})}[(\widehat{\text{BLP}}-\text{BLP}^{*})^{2}]}_{G}+2\underbrace{\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})(\widehat{\text{BLP}}-\text{BLP}^{*})]}_{H}.$
The remainder of the proof shows that $G=\overline{\Delta}^{2}(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$ and $H=0$. Before proceeding, recall that $\overline{\Delta}\coloneqq\sum_{i=1}^{M}\nu_{i}\Delta_{i}/(1+\sum_{i=1}^{M}\nu_{i})$, and note that
$\Sigma^{*-1}\Delta=[\nu_{1}(\Delta_{1}-\overline{\Delta}),\cdots,\nu_{M}(\Delta_{M}-\overline{\Delta}),\mathbf{0}_{N-M}^{\top}]^{\top}$ (18)
obtains from elaborating on (17).
@@
The decomposition exercise in Section 4.3 showed that
@@
Since $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$ by Theorem 2, $\widehat{\Sigma}^{-1}-\Sigma^{*-1}=-\Sigma^{*-1}\Delta\Delta^{\top}\Sigma^{*-1}/(1+\Delta^{\top}\Sigma^{*-1}\Delta)$ by the Sherman-Morrison formula. Thus
@@
and, hence,
$G=\left(\frac{\overline{\Delta}}{1+\Delta^{\top}\Sigma^{*-1}\Delta}\right)^{2}\mathbb{E}_{(b^{*},\Sigma^{*})}\left[(\Delta^{\top}\Sigma^{*-1}(X-b^{*}))^{2}+1+2\Delta^{\top}\Sigma^{*-1}(X-b^{*})\right].$
@@
$=\left(\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\right)^{2}\mathbb{E}[\omega^{2}]+\sum_{i=1}^{M}\nu_{i}^{2}(\Delta_{i}-\overline{\Delta})^{2}\mathbb{E}[\varepsilon_{i}^{2}]$
@@
and inserting this back into the expression for $G$ yields $G=\overline{\Delta}^{2}(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$.
@@
$\omega-\text{BLP}^{*}(X)$ $=\omega-\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b^{*})=\omega-\mathbf{1}_{N}^{\top}\left(\text{diag}(\nu)-\frac{\nu\nu^{\top}}{1+\mathbf{1}_{N}^{\top}\nu}\right)(\omega\mathbf{1}_{N}+\varepsilon)$
$=\frac{\omega-\sum_{i=1}^{N}\nu_{i}\varepsilon_{i}}{1+\mathbf{1}_{N}^{\top}\nu},$
@@
$\Delta^{\top}\Sigma^{*-1}(X-b^{*})$ $=[\nu_{1}(\Delta_{1}-\overline{\Delta}),\cdots,\nu_{M}(\Delta_{M}-\overline{\Delta}),\mathbf{0}_{N-M}^{\top}](\omega\mathbf{1}_{N}+\varepsilon)$
$=\overline{\Delta}\omega+\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\varepsilon_{i}.$
Simplifying $H$ accordingly yields
$H$ $=\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})(\widehat{\text{BLP}}-\text{BLP}^{*})]$
@@
$=\frac{-\overline{\Delta}}{\left(1+\mathbf{1}_{N}^{\top}\nu\right)(1+\Delta^{\top}\Sigma^{*-1}\Delta)}\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\left(\omega-\sum_{i=1}^{N}\nu_{i}\varepsilon_{i}\right)\left(1+\overline{\Delta}\omega+\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\varepsilon_{i}\right)\right]$
$=\frac{-\overline{\Delta}}{(1+\mathbf{1}_{N}^{\top}\nu)(1+\Delta^{\top}\Sigma^{*-1}\Delta)}\left(\overline{\Delta}\mathbb{E}[\omega^{2}]-\sum_{i=1}^{M}\nu_{i}^{2}(\Delta_{i}-\overline{\Delta})\mathbb{E}[\varepsilon_{i}^{2}]\right)$
@@
Proof of Proposition 5.
The solution to the portfolio choice problem presented in Section 5.2 is
@@
under the correctly specified model, and
@@
$\sum_{k=1}^{K}(\alpha^{*}_{k}+\Delta_{\alpha_{k}})\mu_{k}^{*}-\frac{\lambda}{2}\sum_{k=1}^{K}(\alpha^{*}_{k}+\Delta_{\alpha_{k}})^{2}/\eta_{k}-\sum_{k=1}^{K}\alpha^{*}_{k}\mu_{\alpha_{k}}^{*}-\frac{\lambda}{2}\sum_{k=1}^{K}\left(\alpha_{k}^{*}\right)^{2}/\eta_{k}$
@@
and $\sum_{k=1}^{K}\Delta_{\alpha_{k}}=0$. The third equality follows from straightforward algebra.∎

### v3

@@ A.3 Proofs of theorems and propositions
$=-\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\sum_{i=1}^{M}\Omega_{i}^{-i}\Delta_{i}=-\overline{\Delta},$
which completes the proof. ∎
Online Appendix for
“Learning Source Biases: Multisource Misspecifications and Their Impact on Predictions”
by Junnan He, Lin Hu, Matthew Kovach and Anqi Li
Appendix O.1 Further Applications
In this appendix, we utilize the theories developed in the main text to analyze two further applications. The assumption of independent errors is made throughout.
O.1.1 Misperceived media bias
A Republican voter consults several news sources on an evolving issue. Some sources, such as the voter’s personal source and traditional news outlets, carry biases that he feels familiar with. Others, such as social media sites operated by unknown people or bots, are unfamiliar. Initially, the voter is uncertain about the horizontal biases of unfamiliar sources, while holding dogmatic, wrong, beliefs about the biases of some familiar sources. As someone who grew up watching Fox News, he believes that this source is more neutral than it actually is due to familiarity bias (Pennycook and Rand, 2021). He underestimates the bias associated with his own opinions and judgments, in part due to bias blind spot and third-person effect (Pronin et al., 2002; Davison, 1983), [^15] Bias blind spot refers to the tendency for people to recognize biases in human judgment except when that bias is their own. The third-person effect predicts that people perceive mass media messages to have a greater influence on others than on themselves. and in part because he is working toward self-enhancement and would like to view himself as middle-of-the-road (Dennis, 1988). Finally, he believes in a liberal media bias like many other Republicans and displays excessive hostility towards CNN (Hassell et al., 2020). These cognitive biases are known to enhance with age, be resistant to corrective information, or even pass down through generations (Pronin et al., 2002; Jennings et al., 2009). They can be captured by letting ={personal source,Fox $\mathcal{M}=\{\text{personal source},\text{Fox News},\text{CNN}\}$ and $\tilde{b}_{i}<b_{i}^{*}$ $\forall i\in\mathcal{M}$ in our framework.
In each period $t\in\mathbb{N}$, the voter takes a position $a_{t}\in\mathbb{R}$ and incurs a loss $-(a_{t}-\omega_{t})^{2}$ equal to the quadratic distance between this position and the true state of the world. The optimal position in the long run equals $\text{BLP}^{*}(X)$ under the correctly specified model and $\widehat{\text{BLP}}(X)$ under the misspecified model. The expected utilities generated by these positions, evaluated under the true signal distribution, equal $\text{EU}^{*}\coloneqq-\mathbb{E}_{(b^{*},\Sigma^{*})}\left(\omega-\text{BLP}^{*}(X)\right)^{2}$ and $\widehat{\text{EU}}\coloneqq-\mathbb{E}_{(b^{*},\Sigma^{*})}(\omega-\widehat{\text{BLP}}(X))^{2}$, respectively. Their difference captures the objective welfare loss from misspecified learning.
Proposition O.1.
$\widehat{\text{EU}}-\text{EU}^{*}$ equals $-\overline{\Delta}^{2}$ if the covariance matrix of the signals is known to the DM and $-\delta\overline{\Delta}^{2}$ if it is initially unknown.
Proposition O.1, together with Proposition 1 and Theorem 2 of the main text, fully describes the distortionary effects of misspecified learning on the voter’s long-run beliefs, position, and objective welfare. In the current example, $\overline{\Delta}<0$ because $\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}<0$ $\forall i\in\mathcal{M}$. The distortion of the voter’s worldview thus constitutes a liberal media bias of magnitude $|\overline{\Delta}|$. Our results establish a mechanism through which this distortion can spill into the online news sphere and facilitate analysis of its behavioral and welfare consequences.
Hostile media effect and false polarization.
The hostile media effect refers to the phenomenon that opposing partisans perceive news coverage to be more biased against their side than it actually is. Ample evidence in political science, communication, and psychology documents the existence of this phenomenon (Feldman, 2014), [^16] In a classic experiment conducted by Arpan and Raney (2003), student subjects read a balanced story about their hometown college football team in one of three newspapers: home town, rival town, or neutral town outside the state. Subjects viewed the coverage as favoring the rival university, even if it was believed to be from the neutral town. More recently, Hassell et al. (2020) present journalists with balanced news stories that differ only in their ideological content (e.g., the party affiliation of the candidate running for office). Journalists exhibit no gatekeeping bias when selecting news stories to cover, despite the fact that many of them are ideologically liberal, thus debunking the “liberal media bias” that is widely perceived among Republicans. although research on learning and belief formation about new online sources has remained sparse. We demonstrate how the hostile media effect in traditional media can spill into the digital news sphere, as the voter in our model overperceives the bias of any liberal source while underperceiving the bias of any conservative source.
False polarization is defined as the difference between the perceived distance between the positions of two groups and the actual distance between them. Its presence in the American public was rigorously established by Levendusky and Malhotra (2016b). [^17] These authors used survey data from a nationally representative sample that asked about the subjects’ self-placements on an issue scale, as well as their placements of a typical Democratic and Republican voter on that scale. The data revealed that Americans significantly misperceive the public to be more divided along partisan lines than in reality. Since individuals can now express their opinions online at almost no cost, our model predicts a false polarization in the digital news sphere. Furthermore, it suggests a causal relationship between misspecified news bias and false polarization, consistent with the literature’s common attribution of false polarization to partisan media exposure (Levendusky and Malhotra, 2016a).
Fact-checking.
DOOM stipulates that sources with initially unknown biases are irrelevant to the distortionary effect of misspecified learning. As we increase the signal-to-noise ratios of these sources, the voter’s expected utility under the true model and, hence, the misspecified model, increases unambiguously ($\text{EU}^{*}=-(1+\sum_{i\in\mathcal{S}}\nu_{i})^{-1}$ and $\widehat{\text{EU}}=\text{EU}^{*}-{\overline{\Delta}}^{2}$). The same cannot be said about sources with misspecified biases, as increasing their signal-to-noise ratios generally has an ambiguous effect on the welfare distortion $\overline{\Delta}^{2}$.
Our finding suggests that policy interventions that target misinformation and fake news through improving source accuracy (e.g., fact-checking) may entail unintended welfare consequences, especially when they modulate sources that people already hold dogmatic misperceptions about. When it comes to lesser-known sources, fact-checking always improves welfare. Given that professional fact-checking is resource intensive and does not scale easily, the allocation of time and efforts should prioritize lesser-known sources over established ones. This finding is, to our best knowledge, new to the literature on (political) misinformation (see Nyhan 2020 for a survey).
Misperception recalibration.
A large empirical literature studies the treatment effects of misperception recalibration (Bursztyn and Yang, 2022). While early research indicated the possibility of a “backfire effect” whereby recalibration efforts (e.g., provision of corrective information, cross-cutting exposure) end up reinforcing rather than reducing subjects’ misperceptions, an emerging research consensus finds that the backfire effect could be more tenuous than previously thought, and that more attention should be paid to alternative issues, such as the durability and scalability of the interventions (Nyhan, 2021; Hartman et al., 2022). We caution that even if misspecifications can be partially corrected (bring $\Delta_{i}$, $i\in\mathcal{M}$ closer to zero), their behavioral and welfare consequences may remain ambiguous due to the compound learning effect: When source precision is initially unknown, the opportunity to learn about it can render the comparative statics of $\delta\overline{\Delta}$ and $\delta\overline{\Delta}^{2}$ ambiguous.
Discussion.
The economic literature on media bias is vast but is devoted to understanding actual media biases (Anderson et al., eds, 2016). A notable exception is made by Gentzkow et al. (2018) (GWZ), who develop a misspecified learning model of media bias. To model bias, GWZ introduce a payoff-irrelevant confounder, independent of the payoff-relevant state, and call a source biased if it exhibits a nonzero correlation with the confounder. The DM in their model misperceives his personal source as unbiased and uses this belief to guide his media consumption.
The model of GWZ differs from ours in its construct and predictions, in addition to requiring different methodologies to analyze. First, GWZ examine an active learning model where the choice of media consumption is endogenous. Second, they model media bias as a correlation with a random confounder rather than as a horizontal bias, which necessitates the use of very different analytical tools from ours. A consequence of this modeling choice is that (distortions of) the DM’s long-run position has a zero unconditional mean, which contrasts with the prediction of our model. Finally, a central focus of GWZ is on the endogenous trust in news sources, defined as the latter’s (mis)perceived correlation with the true state. Trust is correctly specified throughout our paper and is known to the DM for the most part. When endogenized as a subject of misspecified learning, we find no effect on perceived news bias and only a moderation of the DM’s position due to the compound learning effect.
O.1.2 Project investment under misspecified learning
A manager invests an initial wealth of $W>0$ in $K>1$ projects whose returns $\omega_{k}$, $k=1,\cdots,K$ are independently drawn from the standard normal distribution. Prior to decision making, he inquires about the return of each project from a mix of familiar sources — his personal source, internal consultants — and unfamiliar sources — external consultants, robo-advisors. The signal generated by each source is the same as in the baseline model, with its bias now capturing how optimistic the source is about the project return.
[…]
Over time, the manager observes signal realizations and updates his beliefs about source biases. His long-run beliefs underestimate the biases of sources that are optimistic about project returns, while overestimating the biases of those that are neutral or pessimistic. We examine how such beliefs influence his portfolio choice and objective welfare. For simplicity, we assume that only sources that report on project $k=1$ have their biases misspecified by the manager. The latter has a CARA Bernoulli utility function with an Arrow-Pratt measure $\lambda>0$ of absolute risk aversion.
We work directly with the best linear predictors of asset returns. For each project $k=1,\cdots,K$, we use $\mu_{k}^{*}$ to denote a typical realization of the best linear predictor of its return under the correctly specified model. The counterpart of this quantity under the misspecified model is denoted by $\widehat{\mu}_{k}$. The variance $\eta_{k}^{-1}$ of the best linear predictor is known to decrease with the precision of the signals that report on project $k$. For any given profile of $\mu_{k}$, $k=1,\cdots,K$, the manager solves:
Let $(\alpha_{k}^{*})_{k=1}^{K}$ and $(\widehat{\alpha}_{k})_{k=1}^{K}$ denote the optimal portfolio choices under the correctly specified model and the misspecified model, respectively. The difference $\Delta_{\alpha_{k}}\coloneqq\widehat{\alpha}_{k}-\alpha_{k}^{*}$ captures the distortion in the investment in project $k$. The expected utilities generated by $(\alpha_{k}^{*})_{k=1}^{K}$ and $(\widehat{\alpha}_{k})_{k=1}^{K}$, evaluated under the truth, are denoted by $\text{EU}^{*}$ and $\widehat{\text{EU}}$, respectively. Their difference captures the distortionary effect of misspecified learning on the manager’s objective welfare.
Proposition O.2.
[…]
with decreasing magnitudes in $\lambda$ and $\eta_{k}^{-1}$, $k=1,\cdots,K$.
We perform three comparative static exercises. Consider first the effect of expanding the set of unfamiliar sources, due to the advent of robo-advisors or the popular advice of using “uninvolved managers” to counteract entrenched managerial biases (Guenzel and Malmendier, 2020). This change has no impact on $\overline{\Delta}$ by DOOM while clearly reducing the errors associated with the best linear predictors (as captured by $\eta_{k}^{-1}$, $k=1,\cdots,K$). Consequently, it mitigates the distortionary effects of misspecified learning on the manager’s decision and objective welfare.
Consider next the effect of increasing source precision, which causes $\eta_{k}^{-1}$, $k=1,\cdots,K$ to decrease. Although this change unambiguously mitigates the distortions in the DM’s decision and objective welfare, holding $\overline{\Delta}$ fixed, its impact on $\overline{\Delta}$ is generally ambiguous. This suggests that technological advances that lower data collection and processing costs could impact our manager differently, depending on whether they increase the precision of familiar or novel sources.
Consider next corrective measures aimed at bringing misspecifications closer to zero, which in reality range from de-biasing board members who are prone to select and retain biased CEOs (Guenzel and Malmendier, 2020) to nudges such as premortem, i.e., imagine all twists and turns before decision making (Cohee and Barnhart, 2023). Given that the comparative statics of $\overline{\Delta}$ against these measures are generally ambiguous in the presence of compound misspecified learning, their impact on the manager’s decision and objective welfare can be rather nuanced.
In addition to results that closely mirror those in the main text, our analysis delivers two additional insights. First, suppose we expand the portfolio overseen by the manager. Since this change mechanically increases $K$ and, hence, $\sum_{k=1}^{K}\eta_{k}$, it exacerbates the distortions in the manager’s decision and expected utility (assuming that the signals that report on the newly introduced projects have correctly specified distributions). The intuition is that, in the presence of misspecified learning, adding projects with independent returns to the portfolio creates the illusion that the optimal portfolio is well diversified. Consequently, more funds are allocated to projects the manager is overconfident about, resulting in a greater distortion in the welfare.
Second, all distortionary effects are decreasing in the manager’s degree $\lambda$ of risk aversion.
Appendix O.2 A Frequentist Perspective
In Footnote 3 of the main text, we provided a frequentist foundation for the DM’s limiting belief when the covariance matrix of the signals is known from the outset. When the covariance matrix is also a subject of learning, we claim that the Bayesian asymptotic estimator of $b$ and $\Sigma$ under the normality assumption is equivalent to the joint frequentist estimator without normality assumption that satisfies simultaneously:
  1. 1.
    given $\Sigma$, $b$ is the best least squares linear estimator for means of $X$ (under parametric restriction to $\Theta$), assuming that the true covariance matrix is $\Sigma$;
  2. 2.
    given $b$, $\Sigma$ is the simple sample second moment estimator for covariances of $X$, assuming that the true means are $b$, i.e. $\mathbb{E}[(X-b)(X-b)^{\top}]$.
It is easy to show the first claim: By Footnote 3 of the main text, the estimation of $b$ is obtained by minimizing the population level objective function, assuming that the true sample covariance matrix is $\Sigma$.
To show the second claim, recall the KL divergence formula:
$D_{KL}(b^{*},\Sigma^{*}\|b,\Sigma)=\frac{1}{2}\left(\text{tr}(\Sigma^{-1}\Sigma^{*})-N+(b-b^{*})^{\top}\Sigma^{-1}(b-b^{*})+\log\frac{\text{det}\Sigma}{\text{det}\Sigma^{*}}\right).$
Taking derivative with $\Sigma$ gives the following FOC:
$0=\frac{\partial D_{KL}}{\partial\Sigma}=$ $\frac{1}{2}\left(\frac{\partial}{\partial\Sigma}\text{tr}(\Sigma^{-1}\Sigma^{*})+\frac{\partial}{\partial\Sigma}\text{tr}\left(\Sigma^{-1}(b-b^{*})(b-b^{*})^{\top}\right)+\Sigma^{-1}\right)$
$=$ $\frac{1}{2}\left(-\Sigma^{-1}\Sigma^{*}\Sigma^{-1}-\Sigma^{-1}(b-b^{*})(b-b^{*})^{\top}\Sigma^{-1}+\Sigma^{-1}\right).$
This FOC implies the resulting estimator $\Sigma$ satisfies:
$\Sigma$ $=\Sigma^{*}+(b-b^{*})(b-b^{*})^{\top}$
$=\mathbb{E}[(X-b^{*})(X-b^{*})^{\top}]+(b-b^{*})(b-b^{*})^{\top}$
$=\mathbb{E}[(X-b^{*}+b^{*}-b)(X-b^{*}+b^{*}-b)^{\top}]$
$=\mathbb{E}[(X-b)(X-b)^{\top}],$
where the penultimate line follows from the fact that $X-b^{*}$ is mean zero in expectation (taken under the true parameters $b^{*}$ and $\Sigma^{*}$).
Appendix O.3 Proofs
Proof of Proposition O.1.
In the context of the baseline model, $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\overline{\Delta}$. Consequently,
@@ Discussion.
In the context of Section 5.1 of the main text,
@@
$=-\underbrace{\mathbb{E}_{(b^{*},\Sigma^{*})}[(\widehat{\text{BLP}}-\text{BLP}^{*})^{2}]}_{E}+2\underbrace{\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})(\widehat{\text{BLP}}-\text{BLP}^{*})]}_{F}.$
The proof presented below shows that $E=\overline{\Delta}^{2}(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$ and $F=0$. Before proceeding, recall that $\overline{\Delta}\coloneqq\sum_{i=1}^{M}\nu_{i}\Delta_{i}/(1+\sum_{i=1}^{M}\nu_{i})$, and that
$\Sigma^{*-1}\Delta=\begin{bmatrix}\nu_{1}(\Delta_{1}-\overline{\Delta})\\ \vdots\\ \nu_{M}(\Delta_{M}-\overline{\Delta})\\ \mathbf{0}_{N-M}\end{bmatrix}$
by Equation (3) of the main text.
@@
In Section 5.1 of the main text, we demonstrated that
@@
Since $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$ by Theorem 2, $\widehat{\Sigma}^{-1}=\Sigma^{*-1}-\Sigma^{*-1}\Delta\Delta^{\top}\Sigma^{*-1}/(1+\Delta^{\top}\Sigma^{*-1}\Delta)$ by the Sherman-Morrison formula. Consequently,
@@
and, in turn,
$E=\left(\frac{\overline{\Delta}}{1+\Delta^{\top}\Sigma^{*-1}\Delta}\right)^{2}\mathbb{E}_{(b^{*},\Sigma^{*})}\left[(\Delta^{\top}\Sigma^{*-1}(X-b^{*}))^{2}+1+2\Delta^{\top}\Sigma^{*-1}(X-b^{*})\right].$
@@
$=\left(\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\right)^{2}\mathbb{E}[\omega^{2}]+\sum_{i=1}^{M}\nu_{i}^{2}(\Delta_{i}-\overline{\Delta})^{2}\mathbb{E}[\epsilon_{i}^{2}]$
@@
which, when inserted back into the expression for $E$, yields $E=\overline{\Delta}^{2}(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$.
@@
$\omega-\text{BLP}^{*}(X)$ $=\omega-\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b^{*})=\omega-\mathbf{1}_{N}^{\top}\left(\text{diag}(\nu)-\frac{\nu\nu^{\top}}{1+\mathbf{1}_{N}^{\top}\nu}\right)(\omega\mathbf{1}_{N}+\epsilon)$
$=\frac{\omega-\sum_{i=1}^{N}\nu_{i}\epsilon_{i}}{1+\mathbf{1}_{N}^{\top}\nu},$
@@
$\Delta^{\top}\Sigma^{*-1}(X-b^{*})$ $=[\nu_{1}(\Delta_{1}-\overline{\Delta}),\cdots,\nu_{M}(\Delta_{M}-\overline{\Delta}),\mathbf{0}_{N-M}^{\top}](\omega\mathbf{1}_{N}+\epsilon)$
$=\overline{\Delta}\omega+\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\epsilon_{i}.$
Simplifying $F$ accordingly yields
$F$ $=\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})(\widehat{\text{BLP}}-\text{BLP}^{*})]$
@@
$=\frac{-\overline{\Delta}}{\left(1+\mathbf{1}_{N}^{\top}\nu\right)(1+\Delta^{\top}\Sigma^{*-1}\Delta)}\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\left(\omega-\sum_{i=1}^{N}\nu_{i}\epsilon_{i}\right)\left(1+\overline{\Delta}\omega+\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\epsilon_{i}\right)\right]$
$=\frac{-\overline{\Delta}}{(1+\mathbf{1}_{N}^{\top}\nu)(1+\Delta^{\top}\Sigma^{*-1}\Delta)}\left(\overline{\Delta}\mathbb{E}[\omega^{2}]-\sum_{i=1}^{M}\nu_{i}^{2}(\Delta_{i}-\overline{\Delta})\mathbb{E}[\epsilon_{i}^{2}]\right)$
@@
Proof of Proposition O.2.
The solution to the portfolio choice problem is given by
@@
under the correctly specified model, and it is given by
@@
$\sum_{k=1}^{K}(\alpha^{*}_{k}+\Delta_{\alpha_{k}})\mu_{k}^{*}-\frac{\lambda}{2}\sum_{k=1}^{K}(\alpha^{*}_{k}+\Delta_{\alpha_{k}})^{2}/\eta_{k}-\left[\sum_{k=1}^{K}\alpha^{*}_{k}\mu_{\alpha_{k}}^{*}-\frac{\lambda}{2}\sum_{k=1}^{K}\left(\alpha_{k}^{*}\right)^{2}/\eta_{k}\right]$
@@
and $\sum_{k=1}^{K}\Delta_{\alpha_{k}}=0$ (that is, wealth is fixed), and the third equality follows from straightforward algebra.∎